# CustomContrast: A Multilevel Contrastive Perspective For Subject-Driven Text-to-Image Customization

### v2

@@ Abstract
22footnotetext: Zhendong Mao is the corresponding author.11footnotetext: The project page is CustomContrast.
@@ 1 Introduction
Existing customized T2I methods mainly involve two streams: the finetune-based stream and the finetune-free stream. The finetune-based stream [4, 24, 10] typically fine-tunes each new concept. In this paper, we focus on the finetune-free stream, which achieves real-time generation by training adapters based on the pre-trained encoders. Early work [30, 33] aims to enhance subject similarity by injecting image embeddings into U-Net using additional visual attention. Recent studies further improve text controllability while preserving subject similarity by selectively injecting image features. Some work [12, 37] reduces background interferences by using different reference and reconstruction images, while other work [26] controls the generated regions through additional inference operations. Additionally, some work [27] introduces multimodal large language models to further enhance text controllability. To summarize, existing methods adopt a self-reconstructive perspective, that is, taking target subject images as inputs, aiming to fully mimic all the subject details via the diffusion reconstruction loss [6].
@@ 2.1 Subject-Driven Text-to-image Customization
Existing customized text-to-image work includes finetune-based streams and finetune-free streams. Early work [4, 24, 10, 5, 28, 1, 36, 15] focuses on the finetune-based stream, which requires learning pseudo-words as the subject representations and fine-tuning each new subject. Custom Diffusion [10] and SVDiff [5] reduce computational overhead by minimizing fine-tuning parameters. Recent work [28, 36, 1] in the $\mathcal{P}^{+}$ space uses different pseudo-words for different cross-attention and timesteps.
The finetune-free stream aims to eliminate optimization for each subject, achieving real-time generation. Recent work [30, 12, 7, 37, 26, 17, 29, 18, 27] trains an additional adapter to map reference images to image prompts. Elite [30] uses local mapping to enhance subject features. IP-adapter [33] injects visual features into additional cross-attention, struggling with attribute changes (e.g., shape and orientation). RealCustom [7] and SSR-Encoder [37] selectively inject subject-related features to reduce background features. Other studies [26, 17] use multimodal large language models for image-text alignment. Summarily, existing work follows the self-reconstructive perspective, leading to the trade-off between controllability and similarity.
@@ 2.2 Contrastive Learning of Representations
Early zero-shot classification studies [9, 23] based on CLIP [19] have validated the effectiveness of contrastive learning in aligning visual features with textual features. By dynamically adjusting these representation spaces in downstream tasks, the consistency between visual and textual representations can be further improved. In customized tasks, past work based on the Multimodal Large Language Model [18, 17, 27] has explored using contrastive learning to reduce the gap between image and text spaces. Summarily, similar to traditional contrastive learning, previous studies in customized task primarily focus on aligning image embeddings with text embeddings, usually at the expense of subject similarity. In contrast, our work uses contrastive learning to extract intrinsic features from redundant features. It decouples redundant features through intra-consistency and further learns intrinsic features through inter-distinctiveness, achieving dual optimization of text controllability and subject similarity.
@@ 3.1 Preliminaries
Our work is based on Stable Diffusion (SD) [22], which includes an autoencoder and a UNet denoiser. The encoder $\mathcal{E}(\cdot)$ of the autoencoder maps a given image $\boldsymbol{x}\in\mathbb{R}^{H\times W\times 3}$ to a low-dimensional space $\boldsymbol{z}=\mathcal{E}(\boldsymbol{x})\in\mathbb{R}^{h\times w\times d}$, where $d$ denotes the hidden space dimension. The corresponding decoder $\mathcal{D}(\cdot)$ maps $\boldsymbol{z}$ back to pixel space. In customized task, the text $\boldsymbol{c}_{t}$ and the reference image $\boldsymbol{c}_{i}$ are used as conditional inputs to the denoiser $\epsilon_{\theta}(\cdot)$. Ultimately, the denoiser $\epsilon_{\theta}(\cdot)$ is trained by mean-squared loss:
@@ 3.2 Multimodal Feature Injection Encoder
The CLIP image encoder is used to extract image features. Previous studies [30, 37] show that using only the last layer’s feature of CLIP loses fine-grained details. We propose to concatenate features from different CLIP layers to get coarse-to-fine subject features $\boldsymbol{f}_{I}=\left\{\boldsymbol{f}_{k}\right\}_{k=0}^{K}$, where $K$ refers to layer numbers, defaulting to 3. Then $\boldsymbol{f}_{I}$ is fed into textual and visual Qformer.
Here, Qformer denotes the Transformer module where the queries are learnable queries, and image features are input as part of keys and values. PerceiverAttention (PA) [2] is used in each Qformer to aggregate image features with learnable queries. $PA(\boldsymbol{f}_{I},\boldsymbol{f}_{q})$ aggregates image features $\boldsymbol{f}_{I}$ and learnable queries $\boldsymbol{f}_{q}$ as follows:
@@ 3.3 Multilevel Contrastive Learning Paradigm
3.3.1 Crossmodal Semantic Contrastive Learning
@@
3.3.2 Multiscale Appearance Contrastive Learning
@@ A. $\textit{S}^{+}$ Space
Injecting token $\boldsymbol{s}^{*}\in\mathbb{R}^{k\times d}$ into textual space is used for finetune-based studies [24] as shown in Fig. 3(a). Here, $k$ and $d$ refer to token numbers and dimensions. Define S as the space spanned by $\boldsymbol{s}^{*}$. We then define extended visual concept space $\textit{S}^{+}$ as follows:
@@ B. MACL in $\textit{S}^{+}$ Space
where $\mathcal{D}_{r}(\cdot)$ denotes the ideal distance metric for real subjects $\boldsymbol{x}$, and $\mathcal{D}_{g}(\cdot)$ is relative distance metric for learned features $\boldsymbol{z}$. $i$, $j$ and $k$ denote the $i$-th, $j$-th, and $k$-th samples. We use the CLIP image encoder $\psi_{C}(\cdot)$ to extract appearance features of real subjects, approximating relative distance $\mathcal{D}_{r}$ between real subjects. Note that $\psi_{C}(\cdot)$ is a specific case to approximate $\mathcal{D}_{r}$ and can be replaced by others (e.g., DINO-V2 [25], SigLIP [35]).
@@ 3.4 Timesteps-specific Subject Location
Similar to MOMA[26], we adopt timestep injection inference operations. The main difference is that we use location loss [31] to further reduce redundant features’ interference, as shown in Fig. 4. During training, we extract all visual cross-attention maps $A_{v}$ and p-th text cross-attention maps $A_{t}^{p}$ related to $\textit{S}^{+}$ space (where $p$ is the placeholder index). The global attention map $\overline{A}_{g}$ is then aggregated by averaging these attention maps. $\overline{A}_{g}$ is constrained with the actual mask $M_{g}$ by loss function $\mathcal{L}_{\text{loc}}$ [31]:
@@ 4.1 Experimental Setups
Implementation. Our model is implemented on both SD V1-5 and SDXL. The filtered subset of MVImageNet [34] and OpenImage [11] are used as training sets. The model is trained on 6 A100 GPUs for 200k iterations with learning rate 3e-5 and $\lambda_{1}$=1e-2 and $\lambda_{2}$=1e-3. The layer numbers of Textual and Visual Qformer are set to 4. More details are in Appendix C.1.
@@
Evaluation metrics. We evaluate our model based on DINO-I (DI) [16], CLIP-I (CI), and CLIP-T scores on DreamBench [24]. Similarity. To exclude background interference, we calculate the subject similarity after segmenting the reference subjects and generated subjects with Grounded SAM [21]. Similarity metrics are refined by measuring subject similarity in reconstruction (R-CI and R-DI) and editing (E-CI and E-DI). Controllability. CLIP-T metric is calculated using cosine similarity in the CLIP text-image embedding space. Additionally, ImageReward [32] is also used to evaluate controllability.
@@
CLIP-T $\uparrow$ Image-Reward $\uparrow$ R-CI $\uparrow$ E-CI $\uparrow$ R-DI $\uparrow$ E-DI $\uparrow$
Custom Diffusion 0.258 -1.411 0.863 0.756 0.661 0.517
DreamMatcher 0.298 0.157 0.875 0.781 0.670 0.540
BLIP-Diffusion 0.278 -0.801 0.865 0.768 0.629 0.535
ELITE 0.291 -0.664 0.856 0.762 0.661 0.533
IP-Adapter-plus 0.295 -0.611 0.878 0.766 0.702 0.558
SSR-Encoder 0.303 -0.894 0.857 0.767 0.623 0.524
KOSMOS-G 0.268 -1.446 0.864 0.753 0.677 0.531
MOMA 0.313 -0.243 0.879 0.772 0.679 0.550
ours (SD-v1.5) 0.325 0.454 0.919 0.788 0.737 0.591
@@
Figure 5: Performance visualization with existing SD-v1.5-based methods shows our model’s superior text controllability and subject similarity. The blue line illustrates the dual optimum paradox (controllability and similarity) by adjusting the weight factor of the IP-Adapter [33].
Methods Base-Model controllability similarity
CLIP-T $\uparrow$ Image-Reward $\uparrow$ R-CI $\uparrow$ E-CI $\uparrow$ R-DI $\uparrow$ E-DI $\uparrow$
$\lambda$-ECLIPSE Kv2.2 0.279 -1.031 0.882 0.773 0.677 0.511
IP-Adapter-plus SDXL 0.289 -0.714 0.918 0.790 0.723 0.570
EMU2 SDXL 0.298 -0.203 0.914 0.795 0.719 0.583
MS-Diffusion SDXL 0.312 -0.243 0.901 0.772 0.682 0.540
ours SDXL 0.329 0.446 0.927 0.796 0.742 0.597
Table 2: Quantitative results with more advanced base models. Kv2.2 refers to Kandinsky v2.2 [20].
@@ 4.2 Main Results
Figure 8: Qualitative comparison of human domain generation with existing methods.
Quantitative results. CustomContrast is compared with previous SOTAs on both SD-v1.5 and SDXL. Details of SOTAs are in Appendix C.1. As shown in Fig. 5 and Tab. 1, our model demonstrates significant improvements in both text controllability and image similarity compared to previous SD-V1.5 methods, with a notable increase in the ImageReward metric from 0.157 to 0.454. The quantitative comparison for SDXL, as shown in Tab. 2, indicates that our method also improves both controllability and similarity, with the CLIP-T increasing from 0.312 to 0.329.
To further demonstrate CustomContrast’s capabilities in complex editing, we present qualitative results in multi-object and human domain generation. The qualitative results for the multi-subject task are shown in Fig. 7. Our model can preserve the multiple subject similarities while achieving superior editing capabilities (e.g., only the cat toy wears a hat in the first row; only the cat wears glasses in the third row). To show the generalization ability of CustomContrast, our model is fine-tuned on the human domain dataset FFHQ [8]. As shown in Fig. 8, compared with previous work [13, 31, 33], our model achieves more accurate style transformations (e.g., the second row) and viewpoint changes (e.g., the third row) while preserving ID similarity. This indicates that our model has learned intrinsic ID features and successfully disentangled redundant features.
@@ 4.4 Visualization
To further illustrate our model’s superiority, we use t-SNE to analyze the subject representations of ELITE [30] and our model in Fig. 9. We use dogs with front and side views from DogFaceNet [14] as samples. In Fig. 9(b), the ELITE representations are more entangled by view interference. In contrast, $\boldsymbol{s}^{+}$ learned by CustomContrast better preserves relative distances between samples in Fig. 9(c), with different views of the same dog being closer. These analyses demonstrate our model decouples view redundant features from intrinsic subject features.
@@ 5 Conclusion
Appendix A Methodology supplement
A.1 Multiscale Appearance Contrastive Learning
First, let’s review the objectives that MACL aims to achieve. MACL in the $\mathcal{S}^{+}$ space is designed to preserve the multi-scale similarity structure, meaning that the similarity of learned features should be positively correlated with the similarity of real subjects. Consequently, the similarity between samples must satisfy the following conditions:
where $\mathcal{D}_{r}(\cdot)$ represents the ideal distance metric for reference subjects, while $\mathcal{D}_{g}(\cdot)$ denotes the distance metric for the subject representation. The subscripts $i$, $j$ and $k$ denote the $i$-th, $j$-th, and $k$-th samples, respectively.
In the main text, we primarily achieve relative distance preservation with the real subjects through the scaling factor. To further explain the role of the scaling factor, we derive the MACL loss gradient corresponding to the $a$-th cross-attention layer with negative similarity:
$\frac{\partial\mathcal{L}_{m}^{ij}}{\partial cos_{ij}^{a}}$ $=\frac{1}{\tau r_{ij}}\frac{\exp\left[cos_{ij}^{a}/\left(\tau r_{ij}\right)\right]}{\sum_{k\neq i}\exp\left[cos_{ik}^{a}/\left(\tau r_{ik}\right)\right]},$ (11)
$cos_{ij}^{a}$ $\leftarrow cos_{ij}^{a}-\zeta\frac{\partial\mathcal{L}_{m}^{ij}}{\partial cos_{ij}^{a}},$
where $\zeta$ is the learning rate and $r_{ij}$ is the cosine similarity between the i-th real subject and the j-th real subject. As indicated by Eq. 11, the gradient of $cos_{ij}^{a}$ is inversely proportional to $r_{ij}$. A larger similarity $r_{ij}$ corresponds to a smaller gradient. That is, for samples with low real subject similarities, MACL will push it further with a stronger amplitude. In contrast, for samples with high real subject similarities, MACL will push it away with a relatively weak amplitude. Ultimately, the condition of Eq. 10 could be approximated, preserving the real subject similarity structure.
Appendix B Training Strategies
B.1 Multi-perspective contrastive learning
Figure 10: Data processing of multi-view images for contrastive learning. Given a set of multi-view images, we start by sampling three frames and segmenting the subject in each. Then the subject of the first frame is injected into the image encoder, the middle frame serves as the reference subject for contrastive learning, and the last frame is used as reconstruction supervision.
Previous work [30, 10, 3] has applied rotation, cropping, and affine transformation to a single image, or utilized different frames from a video as reference images and reconstruction supervision. However, these methods fall short in fully capturing the dynamic nature of the physical world. To alleviate this problem, we propose a data sampling method for multi-view contrastive learning.
Figure 11: Qualitative results with existing methods on SD-v1.5. Compared to previous models based on SD-v1.5, our method excels in text editability and individual similarity. It performs well even with complex shape (e.g., turning a vase into a cup in the second row) and accessorization modification (e.g., clock with angle wings in the fourth row), where other models underperform.
The data preparation process is illustrated in Fig. 10. For a set of multi-view images in MVImageNet [34], we sample three frames and segment the subject. For the first frame, the segmented foreground subject serves as the reference image, and the subject of the middle frame is also segmented out as a positive example of contrastive learning. The cropped last frame is used as reconstruction supervision. The intermediate frame used for contrastive learning serves as a bridge, enhancing the alignment between the first frame and the final frame. Additionally, it enables the model to learn the features of the subject accurately while excluding background interference.
Method Finetune Free Base Model Freezing MLLM Free
Custom Diffusion
DreamMatcher
ELITE
BLIP-Diffusion
IP-Adapter
SSR-Encoder
MOMA
Kosmos-G
$\lambda$-ECLIPSE
Emu2
MS-Diffusion
Ours
Table 5: An overview of compared previous SOTAs on the subject-driven text-to-image customization.
In the multi-view MVImageNet dataset, contrastive learning frames naturally contain different views’ features compared to the reference images. During training, we crop different sizes centered on the subject to simulate different subject sizes randomly. Additionally, we apply random padding horizontally or vertically to alter the subject’s position. Since the OpenImage dataset [11] does not contain multi-view sample pairs, we use some data-augmented images (e.g., resizing, rotation, and cropping) as positive samples in this dataset.
Appendix C More Experiments Results
C.1 implementation details
We use filtered subset of OpenImage [11] and MVImageNet [34] trainset. On these datasets, we discard samples that are related to people or where the subject is too large (0.85) or too small (0.15) relative to the image area. This results in a 2M dataset for training. These datasets inherently have segmentation maps. Additionally, we use BLIP-2 to generate captions for the reference images and employ a NER model to extract entities from these captions.
Style Modification Shape Modification
A watercolor painting of {} on a mountain a {} in the shape of a bag
A painting of {} in Van Gogh style a {} in the shape of a cat
{} in the style of Banksy art a {} in the shape of a dog
an oil painting of a {} a {} in the shape of a cup
Scene Modification Accessorization Modification
a {} riding a bike in the city a {} wearing a hat on the beach
a {} floating in a pond a {} wearing sunglasses in a garden
a {} in a schoolyard playground a {} has angle wings on the sky
a {} on a sandy beach near the dunes a {} wearing purple wizard outfit and playing guitar On the stage
a {} on a wooden dock by a lake a {} wearing bowtie and top hat in the office
a {} at the edge of a swimming pool a {} wearing a scarf in the snow
Table 6: text prompt list. Evaluation prompts in complex application scenarios. ”{}” represents the class of the corresponding subject.
The layer numbers of Textual Qformer and Visual Qformer are set to 4, while the TV Fusion Module layer number is set to 1. The visual query number is set to 17. The textual token of textual Qformer is set to 17 in SD-v1.5 and 71 in SDXL. The extra one is CLS token, which is used for Crossmodal Semantic Contrastive Learning. The DDIM sampler is used for inference and the guidance scale is set to 7.5. During inference, $\gamma$ is set to 0.55.
Prior SOTAs. We compare CustomContrast with open source state-of-the-arts, divided into two categories: finetune-based work: Custom Diffusion [10], DreamMatcher [15]; finetune-free work: ELITE [30], BLIP-Diffusion [12], IP-Adapter [33], $\lambda$-ECLIPSE [18], SSR-Encoder [37], MS-Diffusion [29], MOMA [26], KOSMOS-G [17], and EMU2 [27]. State-of-the-art work without public code is not compared. As shown in Tab. 5, we present the specific differences in previous studies. ”MLLM Free” indicates no reliance on the Multimodal Large Language Model.
During inference, we use the default settings of each method. Each method generates 16 images for each subject in DreamBench [24] based on each specified test prompt. Note that DreamMatcher [15] is based on Custom Diffusion [10]. When generating multiple subjects, EMU2[27] and MS-Diffusion[29] both use layout control. To maintain consistency, we use the same layout area for these two models. At the same time, our model also uses the same area of layout control when generating multiple subjects.
C.2 Test Prompt List
we create a prompt set divided into four categories: style modification, shape modification (such as animal species), scene modification, and accessorization modification. The specific prompt set is shown in Tab. 6.
Figure 12: Qualitative comparison between CustomContrast and existing methods based on SDXL. When faced with complex shape modifications (e.g., the first row) and accessorization modifications (e.g., the fourth row), CustomContrast still performs excellently.
C.3 More qualitative results based on SD-v1.5.
As shown in Fig. 11, we present further qualitative comparisons between CustomContrast trained on SD-v1.5 and other SD-v1.5-based models. Attributes within an image are complexly coupled, leading previous methods to misconstrue redundant attributes as the subject’s intrinsic properties. During inference, these redundant features limit the flexible generation of the subject. For example, in the first row, IP-Adapter[33] and SSR-Encoder[37] retain the original image’s background, which is unnecessary for scene modification; in the second row, redundant features learned from the vase’s shape make it difficult for previous methods to modify the vase into a cup. CustomContrast adopts contrastive learning to extract intrinsic features from redundant features, better achieving various complex personalized generations.
C.4 More qualitative results based on SDXL.
As shown in Fig. 12, we present further qualitative comparisons between CustomContrast trained on SDXL and other SDXL-based models.
Figure 13: Visualization for timesteps-specific subject location on SD-v1.5. As the timesteps increase, the global attention maps $\overline{A}_{g}$ converge from a scattered state to the subject itself.
[…]
0 1 0.774 0.574 0.323
1 2 0.781 0.580 0.321
2 3 0.788 0.591 0.325
3 5 0.789 0.596 0.317
Table 7: Ablation studies on the effectiveness of CLIP-Image Encoder layers. When using three intermediate layers of the CLIP Image Encoder, both subject similarity and text controllability reach a good balance.
C.5 More Ablation Studies
In this section, we will present more ablation experiments. All ablation experiments are trained on SD-v1.5.
Effectiveness of layer numbers used by CLIP Image-encoder. Previous studies [30, 37, 28] have pointed out that using only the last layer features of the CLIP image encoder tends to lose detailed appearance features. To better capture the fine details of the reference image, we extract features from different intermediate layers of CLIP to retain more fine-grained details. As shown in Tab. 7, using only a single layer of CLIP features results in low subject similarity. As the number of layers increases, subject similarity improves, but often at the expense of text controllability. When using three intermediate layers of the CLIP Image Encoder, both subject similarity and text controllability reach a good balance. Therefore, we ultimately use 3 intermediate layers.
C.6 Visualization for Timesteps-specific Subject Location
Specifically, we use bilinear interpolation to resize different cross-attention maps to a size of $64\times 64$ and then average them to aggregate the global attention map $\overline{A}_{g}$. As shown in Fig. 13, as the timestep increases, the extracted global attention $\overline{A}_{g}$ gradually focuses on the subject itself. That is, in the denoising process, at a smaller timestep, the subject’s global attention map $\overline{A}_{g}$ is scattered throughout the entire image, while at a larger timestep, the subject’s global attention map $\overline{A}_{g}$ will gradually focus on the subject itself (e.g., in the first row, the clock’s attention is almost entirely focused on itself at the 50-th timesteps). This timestep injection operation helps the model more accurately inject the subject’s intrinsic features while reducing the injection of redundant features.

### v3

@@ Abstract
codehttps://cn-makers.github.io/CustomContrast/
@@ 1 Introduction
Existing customized T2I methods mainly involve two streams: the finetune-based stream and the finetune-free stream. The finetune-based stream (Gal et al. 2022; Ruiz et al. 2023; Kumari et al. 2023) typically fine-tunes each new concept. In this paper, we focus on the finetune-free stream, which achieves real-time generation by training adapters based on the pre-trained encoders. Early work (Wei et al. 2023; Ye et al. 2023) aims to enhance subject similarity by injecting image embeddings into U-Net using additional visual attention. Recent studies further improve text controllability while preserving subject similarity by selectively injecting image features. Some work (Li, Li, and Hoi 2024; Zhang et al. 2024) reduces background interferences by using different reference and reconstruction images, while other work (Song et al. 2024) controls the generated regions through additional inference operations. Additionally, some work (Sun et al. 2024) introduces multimodal large language models to further enhance text controllability. To summarize, existing methods adopt a self-reconstructive perspective, that is, taking target subject images as inputs, aiming to fully mimic all the subject details via the diffusion reconstruction loss (Ho, Jain, and Abbeel 2020).
@@ 2.1 Subject-Driven Text-to-image Customization
Existing customized text-to-image work includes finetune-based streams and finetune-free streams. Early work (Gal et al. 2022; Ruiz et al. 2023; Kumari et al. 2023; Han et al. 2023; Voynov et al. 2023; Alaluf et al. 2023; Zhang et al. 2023; Nam et al. 2024; Huang et al. 2022) focuses on the finetune-based stream, which requires learning pseudo-words as the subject representations and fine-tuning each new subject. Custom Diffusion (Kumari et al. 2023) and SVDiff (Han et al. 2023) reduce computational overhead by minimizing fine-tuning parameters. Recent work (Voynov et al. 2023; Zhang et al. 2023) in the $\mathcal{P}^{+}$ space uses different pseudo-words for different cross-attention and timesteps.
The finetune-free stream aims to eliminate optimization for each subject, achieving real-time generation. Recent work (Wei et al. 2023; Li, Li, and Hoi 2024; Huang et al. 2024; Zhang et al. 2024; Song et al. 2024; Pan et al. 2023; Wang et al. 2024; Patel et al. 2024; Sun et al. 2024; Chen et al. 2023; Huang et al. 2023b, a) trains an additional adapter to map reference images to image prompts. Elite (Wei et al. 2023) uses local mapping to enhance subject features. IP-adapter (Ye et al. 2023) injects visual features into additional cross-attention, struggling with attribute changes (e.g., shape and orientation). RealCustom (Huang et al. 2024; Mao et al. 2024) and SSR-Encoder (Zhang et al. 2024) selectively inject subject-related features to reduce background features. Other studies (Song et al. 2024; Pan et al. 2023) use multimodal large language models for image-text alignment.
@@ 2.2 Contrastive Learning of Representations
Early zero-shot classification studies (Khattak et al. 2023; Roy and Etemad 2024) based on CLIP (Radford et al. 2021) have validated the effectiveness of contrastive learning in aligning visual features with textual features. By dynamically adjusting these representation spaces in downstream tasks, the consistency between visual and textual representations can be further improved. In customized tasks, past work based on the Multimodal Large Language Model (MLLM) (Patel et al. 2024; Pan et al. 2023; Sun et al. 2024) has explored using contrastive learning to reduce the gap between image and text spaces. Summarily, similar to traditional contrastive learning, previous studies in customized task primarily focus on aligning image embeddings with text embeddings, usually at the expense of subject similarity. In contrast, our work uses contrastive learning to extract intrinsic features from redundant features. It decouples redundant features through intra-consistency and further learns intrinsic features through inter-distinctiveness, achieving dual optimization of text controllability and subject similarity.
@@ 3.1 Preliminaries
Our work is based on Stable Diffusion (SD) (Rombach et al. 2022), which includes an autoencoder and a UNet denoiser. The encoder $\mathcal{E}(\cdot)$ of the autoencoder maps a given image $\boldsymbol{x}\in\mathbb{R}^{H\times W\times 3}$ to a low-dimensional space $\boldsymbol{z}=\mathcal{E}(\boldsymbol{x})\in\mathbb{R}^{h\times w\times d}$, where $d$ denotes the hidden space dimension. The corresponding decoder $\mathcal{D}(\cdot)$ maps $\boldsymbol{z}$ back to pixel space. In customized task, the text $\boldsymbol{c}_{t}$ and the reference image $\boldsymbol{c}_{i}$ are used as conditional inputs to the denoiser $\epsilon_{\theta}(\cdot)$. Ultimately, the denoiser $\epsilon_{\theta}(\cdot)$ is trained by mean-squared loss:
@@ 3.2 Multimodal Feature Injection Encoder
The CLIP image encoder is used to extract image features. Previous studies (Wei et al. 2023; Zhang et al. 2024) show that using only the last layer’s feature of CLIP loses fine-grained details. We propose to concatenate features from different CLIP layers to get coarse-to-fine subject features $\boldsymbol{f}_{I}=\left\{\boldsymbol{f}_{k}\right\}_{k=0}^{K}$, where $K$ refers to layer numbers, defaulting to 3. Then $\boldsymbol{f}_{I}$ is fed into textual and visual Qformer.
Here, Qformer denotes the Transformer module where the queries are learnable queries, and image features are input as part of keys and values. PerceiverAttention (PA) (Awadalla et al. 2023) is used in each Qformer to aggregate image features with learnable queries. $PA(\boldsymbol{f}_{I},\boldsymbol{f}_{q})$ aggregates image features $\boldsymbol{f}_{I}$ and learnable queries $\boldsymbol{f}_{q}$ as follows:
@@ 3.3 Multilevel Contrastive Learning Paradigm
Crossmodal Semantic Contrastive Learning
@@
Multiscale Appearance Contrastive Learning
@@ A. $\textit{S}^{+}$ Space
Injecting token $\boldsymbol{s}^{*}\in\mathbb{R}^{k\times d}$ into textual space is used for finetune-based studies (Ruiz et al. 2023) as shown in Fig. 3(a). Here, $k$ and $d$ refer to token numbers and dimensions. Define S as the space spanned by $\boldsymbol{s}^{*}$. We then define extended visual concept space $\textit{S}^{+}$ as follows:
@@ B. MACL in $\textit{S}^{+}$ Space
where $\mathcal{D}_{r}(\cdot)$ denotes the ideal distance metric for real subjects $\boldsymbol{x}$, and $\mathcal{D}_{g}(\cdot)$ is relative distance metric for learned features $\boldsymbol{z}$. $i$, $j$ and $k$ denote the $i$-th, $j$-th, and $k$-th samples. We use the CLIP image encoder $\psi_{C}(\cdot)$ to extract appearance features of real subjects, approximating relative distance $\mathcal{D}_{r}$ between real subjects. Note that $\psi_{C}(\cdot)$ is a specific case to approximate $\mathcal{D}_{r}$ and can be replaced by others (e.g., DINO-V2 (Sinhamahapatra et al. 2024), SigLIP (Zhai et al. 2023)).
@@ 3.4 Timesteps-specific Subject Location
Similar to MOMA(Song et al. 2024), we adopt timestep injection inference operations. The main difference is that we use location loss (Xiao et al. 2023) to further reduce redundant features’ interference, as shown in Fig. 4. During training, we extract all visual cross-attention maps $A_{v}$ and p-th text cross-attention maps $A_{t}^{p}$ related to $\mathcal{S}^{+}$ space (where $p$ is the placeholder index). The global attention map $\overline{A}_{g}$ is then aggregated by averaging these attention maps. $\overline{A}_{g}$ is constrained with the actual mask $M_{g}$ by loss function $\mathcal{L}_{\text{loc}}$ (Xiao et al. 2023):
@@ 4.1 Experimental Setups
Implementation. Our model is implemented on both SD V1-5 and SDXL. The filtered subset of MVImageNet (Yu et al. 2023) and OpenImage (Kuznetsova et al. 2020) are used as training sets. The model is trained on 6 A100 GPUs for 200k iterations with learning rate 3e-5 and $\lambda_{1}$=1e-2 and $\lambda_{2}$=1e-3. The layer numbers of Textual and Visual Qformer are set to 4. More details are in Appendix C.1.
@@
Evaluation metrics. We evaluate our model based on DINO-I (DI) (Oquab et al. 2023), CLIP-I (CI), and CLIP-T (CT) scores on DreamBench (Ruiz et al. 2023). Similarity. To exclude background interference, we calculate the subject similarity after segmenting the reference subjects and generated subjects with Grounded SAM (Ren et al. 2024). Similarity metrics are refined by measuring subject similarity in reconstruction (R-CI and R-DI) and editing (E-CI and E-DI). Controllability. CLIP-T metric is calculated using cosine similarity in the CLIP text-image embedding space. Additionally, ImageReward (IR) (Xu et al. 2024) is also used to evaluate controllability.
@@
CT $\uparrow$ IR $\uparrow$ R-CI $\uparrow$ E-CI $\uparrow$ R-DI $\uparrow$ E-DI $\uparrow$
Custom Diffusion 0.258 -1.41 0.863 0.756 0.661 0.517
DreamMatcher 0.298 0.16 0.875 0.781 0.670 0.540
BLIP-Diffusion 0.278 -0.80 0.865 0.768 0.629 0.535
ELITE 0.291 -0.66 0.856 0.762 0.661 0.533
IP-Adapter-plus 0.295 -0.61 0.878 0.766 0.702 0.558
SSR-Encoder 0.303 -0.89 0.857 0.767 0.623 0.524
KOSMOS-G 0.268 -1.45 0.864 0.753 0.677 0.531
MOMA 0.313 -0.24 0.879 0.772 0.679 0.550
ours (SD-v1.5) 0.325 0.45 0.919 0.788 0.737 0.591
@@
Methods controllability similarity
CT $\uparrow$ IR $\uparrow$ R-CI $\uparrow$ E-CI $\uparrow$ R-DI $\uparrow$ E-DI $\uparrow$
$\lambda$-ECLIPSE 0.279 -1.03 0.882 0.773 0.677 0.511
IP-Adapter-XL 0.289 -0.71 0.918 0.790 0.723 0.570
EMU2 0.298 -0.203 0.91 0.795 0.719 0.583
MS-Diffusion 0.312 -0.24 0.901 0.772 0.682 0.540
ours (SDXL) 0.329 0.45 0.927 0.796 0.742 0.597
Table 2: Quantitative results with more advanced base models, i.e., SDXL and kv2.2 (Razzhigaev et al. 2023).
@@ 4.2 Main Results
Quantitative results. CustomContrast is compared with previous SOTAs on both SD-v1.5 and SDXL. Details of SOTAs are in Appendix C.1. As shown in Tab. 1, our model demonstrates significant improvements in both text controllability and image similarity compared to previous SD-V1.5 methods, with a notable increase in the ImageReward metric from 0.157 to 0.454. The quantitative comparison for SDXL, as shown in Tab. 2, indicates that our method also improves both controllability and similarity, with the CLIP-T increasing from 0.312 to 0.329.
@@
To further demonstrate CustomContrast’s capabilities in complex editing, we present qualitative results in multi-object and human domain generation. The qualitative results for the multi-subject task are shown in Fig. 6. Our model can preserve the multiple subject similarities while achieving superior editing capabilities (e.g., only the cat toy wears a hat in the first row; only the cat wears glasses in the third row). Human domain qualitative results are in Appendix C.5.
@@ 4.4 Visualization
To further illustrate our model’s superiority, we use t-SNE to analyze the subject representations of ELITE (Wei et al. 2023) and our model in Fig. 7. We use dogs with front and side views from DogFaceNet (Mougeot, Li, and Jia 2019) as samples. Different shapes represent different dogs, and different colors indicate different views. In Fig. 7(b), the ELITE representations are more entangled by view interference. In contrast, $\boldsymbol{s}^{+}$ learned by CustomContrast better preserves relative distances between samples in Fig. 7(c), with different views of the same dog being closer. These analyses demonstrate our model decouples view redundant features from intrinsic subject features.
@@ 5 Conclusion
Acknowledgements
This research is supported by Artificial Intelligence-National Science and Technology Major Project 2023ZD0121200, National Natural Science Foundation of China under Grant 62222212 and National Natural Science Foundation of China under Grant 623B2094.